If both mass and radius of earth shrinks by 50% then acceln due to gravity changes by

Grade:12

1 Answers

DR STRANGE
7 years ago
 
Hi Nizam,
g=GM/R2
g’=G(M/2)/(R/2)2=2GM/R2=2g..................................................ans
Approve if you are satisfied :)